### 26 Feb 2024 v2

# Cylindrical Gravitational Waves in Einstein-Aether Theory

Abstract
Along the lines of the Einstein-Rosen wave equation of General Relativity (GR), we derive a gravitational wave equation with cylindrical symmetry in the Einstein-aether (EA) theory. We show that the gravitational wave in the EA is periodic in time for both the metric functions $\Psi(r,t)$ and $H(r,t)$. However, in GR, $\Psi(r,t)$ is periodic in time, but $H(r,t)$ is semi-periodic in time, having a secular drifting in the wave frequency. The evolution of wave pulses of a given width is entirely different in both theories in the $H(r,t)$ metric function due to this frequency drifting. Another fundamental difference between the two theories is the gravitational wave velocity. While in GR, the waves propagate with the speed of light, in EA, there is no upper limit to the wave velocity, reaching infinity if $c_{13}\rightarrow 1$ and zero if $c_{13}\rightarrow-\infty$. We also show that energy-momentum pseudotensor and superpotential get contributions from aether in addition to the usual gravitational field part. All these characteristics are observational signatures that differentiate GR and EA.
1 Introduction
Immediately after completing the theory of General Relativity [1] and proposing the three classical tests [2] to verify the theory experimentally, in 1916, Einstein predicted the existence of gravitational waves as a natural consequence of dynamical spacetime [3]. In this first article [3] on gravitational waves, Einstein concluded that accelerated massive objects generate changes in the spacetime curvature, which propagate like waves with the speed of light. However, he concluded that gravitational waves carry no energy because of a calculational error, which he corrected in 1918 [4]. He also derived a famous formula for the energy loss of a system emitting gravitational radiation. Nevertheless, Einstein tried to disprove the existence of gravitational waves in 1937 [5]. The issue of gravitational waves remained controversial until Bondi [6] settled it for good, which was later explored in a series of papers by Bondi and his collaborators. The first experimental evidence, albeit indirect, came in 1979 with the discovery of the orbital decay of the newly discovered binary pulsar by Hulse and Taylor [7]. Finally, the gravitational waves were directly detected in 2015 by the LIGO and Virgo collaborations [8]. Now, with this direct detection of gravitational waves, we have a new window into the universe and can test our understanding of gravity at higher energies like never before. A fascinating account of the history of gravitational waves can be found in references [9, 10, 11].
Einstein and Rosen were the first to consider cylindrical gravitational waves. They solved the field equations of the time-dependent vacuum with cylindrical symmetry [5]. This solution described the outer spacetime for a collapsing cylindrical source, producing gravitational waves. In fact, cylindrical symmetry is the simplest symmetry capable of producing gravitational waves, as the Birkhoff theorem forbids the production of gravitational waves from spherical symmetry [12]. A monochromatic wave, or a pulse of cylindrical symmetry, moves inward in matter-free space, implodes on the axis, and moves out again. This is the only problem of gravitational radiation where one has an exact solution to the general relativity field equations. Although this special solution does not illustrate all the salient features of gravitational radiation, we can still make correct general statements about gravitational waves. This problem, therefore, occupies a unique position in the theory of gravitational radiation. Also, the cylindrical spacetimes are of great interest, as they allow the study of a wide range of physical phenomena, from the gravitational analog of the Aharonov-Bohm effect to the cosmic superstrings. A recent review of this topic can be found in [13]. The Einstein-Rosen waves are still investigated from very different perspectives, from the Renormalization group approach [14] to the Geroch group [15]; from chaos theory [16] to holography [17]; and from gravitational collapse [18] to observable quantum gravity effects [19]. More on cylindrical gravitational waves can be found in the recent articles [20] and [21].
Since the first observation of gravitational waves, several works have studied the gravitational waves in EA theory, for example, the references [22], [23], [24] and [25]. Even before that, Jacobson and Mattingly [26] obtained the gravitational wave modes by linearizing the EA field equations. In a previous work, Chan & da Silva (2021) [27] present all the possible solutions for a static cylindrical symmetric spacetime in the EA theory. In this work, we obtain the analog of the Einstein-Rosen wave equations using the full EA field equations without the linearization.
The paper is organized as follows. In Section 2, we present the EA field equations. In Section 3, we compute field equations for a general cylindrical metric, obtain the cylindrical gravitational wave equation in Section 4, and solve it in Section 5. In Sections 6 and 7, we study the gravitational wave pulse. We compute energy-momentum pseudotensor and superpotential in Section 8. We summarize our results in Section 9.
2 Field equations in EA theory
Here, we give a very brief summary of the EA theory. We refer the reader to our earlier papers on the subject for more details [29, 30, 31, 32, 28]. The general action of the EA theory, in a background where the metric signature is (+++)({-}{+}{+}{+})( - + + + ) and with the units chosen such that the speed of light defined by the metric $g_{ab}$ is unity, is given by
$S=\int\sqrt{-g}~{}(L_{\rm Einstein}+L_{\rm aether}+L_{\rm matter})d^{4}x,$ (1)
where the first term is the usual Einstein-Hilbert Lagrangian, defined by $R$, the Ricci scalar, and $G$, the EA gravitational constant, as
$L_{\rm Einstein}=\frac{1}{16\pi G}R.$ (2)
The second term, the aether Lagrangian, is given by
$L_{\rm aether}=\frac{1}{16\pi G}[-K^{ab}{}_{mn}\nabla_{a}u^{m}\nabla_{b}u^{n}+\lambda(g_{ab}u^{a}u^{b}+1)],$ (3)
where the tensor ${K^{ab}}_{mn}$ is defined as
${K^{ab}}_{mn}=c_{1}g^{ab}g_{mn}+c_{2}\delta^{a}_{m}\delta^{b}_{n}+c_{3}\delta^{a}_{n}\delta^{b}_{m}-c_{4}u^{a}u^{b}g_{mn},$ (4)
being the $c_{i}$ dimensionless coupling constants, and $\lambda$ a Lagrange multiplier enforcing the unit timelike constraint on the aether, and
$\delta^{a}_{m}\delta^{b}_{n}=g^{a\alpha}g_{\alpha m}g^{b\beta}g_{\beta n}.$ (5)
The last term, $L_{\rm matter}$, is the matter Lagrangian, which depends on the metric tensor and the matter field.
In the weak-field, slow-motion limit EA theory reduces to Newtonian gravity with a value of Newton’s constant $G_{\rm N}$ related to the parameter $G$ in the action (1) by [33],
$G=G_{N}\left(1-\frac{c_{14}}{2}\right).$ (6)
Here, the constant $c_{14}$ is defined as
$c_{14}=c_{1}+c_{4}.$ (7)
The field equations are obtained by extremizing the action with respect to independent variables of the system. The variation with respect to the Lagrange multiplier $\lambda$ imposes the condition that $u^{a}$ is a unit timelike vector, thus
$g_{ab}u^{a}u^{b}=-1,$ (8)
while the variation of the action with respect $u^{a}$, leads to [33]
$\nabla_{a}J^{a}_{b}+c_{4}u^{b}\nabla_{b}u_{a}\nabla_{b}u^{a}+\lambda u_{b}=0,$ (9)
where,
$J^{a}_{m}=K^{ab}_{mn}\nabla_{b}u^{n}.$ (10)
The variation of the action with respect to the metric $g_{mn}$ gives the dynamical equations,
$G^{Einstein}_{ab}=T^{aether}_{ab}+8\pi GT^{matter}_{ab},$ (11)
where
$G^{Einstein}_{ab}$ $=$ $R_{ab}-\frac{1}{2}g_{ab}R,$
$T^{aether}_{ab}$ $=$ $\nabla_{c}[J^{c}\;_{(a}u_{b)}+u^{c}J_{(ab)}-J_{(a}\;^{c}u_{b)}]$
$-\frac{1}{2}g_{ab}J^{c}\;_{d}\nabla_{c}u^{d}+\lambda u_{a}u_{b}$
$+c_{1}[\nabla_{a}u_{c}\nabla_{b}u^{c}-\nabla^{c}u_{a}\nabla_{c}u_{b}]$
$+c_{4}u^{c}\nabla_{c}u_{a}u^{c}\nabla_{c}u_{b},$
$G^{EA}_{ab}$ $=$ $G^{Einstein}_{ab}-T^{aether}_{ab},$
$T^{matter}_{ab}$ $=$ $\frac{-2}{\sqrt{-g}}\frac{\delta\left(\sqrt{-g}L_{matter}\right)}{\delta g_{ab}}.$ (12)
In a more general situation, the Lagrangian of GR is recovered if and only if the coupling constants are identically zero, e.g., $c_{1}=c_{2}=c_{3}=c_{4}=0$, considering the equations (4) and (8).
3 Cylindrical gravitational waves
Let us start with the most general static cylindrical symmetric, given by
$ds^{2}$ $=$ $-e^{2H(r,t)-2\Psi(r,t)}dt^{2}+e^{2H(r,t)-2\Psi(r,t)}dr^{2}$ (13)
$+e^{-2\Psi(r,t)}r^{2}d\varphi^{2}+e^{2\Psi(r,t)}dz^{2}.$
We assume that the indices of the Riemann e Einstein tensors as $(0,1,2,3)$ corresponding to the coordinates $(t,r,\varphi,z)$, respectively. In accordance with equation (8), the aether field is assumed unitary, timelike, and constant, chosen as
$u^{a}=(e^{\Psi(r,t)-H(r,t)},0,0,0).$ (14)
Assuming (13), we compute the different terms in the field equations (11), giving
$(-2H_{1}\Psi_{1}r-2rH_{0}\Psi_{0}-2H_{1}+3\Psi_{0}^{2}r+2\Psi_{1}$
$+\Psi_{1}^{2}r-2rH_{11}+rH_{0}^{2}+2\Psi_{11}r+rH_{1}^{2})c_{13}$
$+(\Psi_{0}^{2}r-2rH_{0}\Psi_{0}+rH_{0}^{2})c_{2}$
$+(2rH_{11}-2\Psi_{1}-2\Psi_{11}r-\Psi_{1}^{2}r+2H_{1}$
$+2H_{1}\Psi_{1}r-rH_{1}^{2})c_{3}+(-2H_{1}-2H_{1}\Psi_{1}r$
$+rH_{1}^{2}-2rH_{11}+2\Psi_{1}+\Psi_{1}^{2}r+2\Psi_{11}r)c_{4}$
$+2H_{1}-2\Psi_{1}^{2}r-2\Psi_{0}^{2}r=0,$ (15)
$(\Psi_{0}H_{1}r-3\Psi_{1}\Psi_{0}r-H_{1}H_{0}r+H_{0}+H_{0}\Psi_{1}r)c_{13}$
$+(H_{0}\Psi_{1}r-\Psi_{1}\Psi_{0}r-H_{1}H_{0}r+\Psi_{0}H_{1}r)c_{2}$
$-H_{0}+2\Psi_{1}\Psi_{0}r=0,$ (16)
$(H_{0}\Psi_{1}-\Psi_{1}\Psi_{0}-H_{1}H_{0}+\Psi_{0}H_{1})rc_{14}-$
$H_{0}+2\Psi_{1}\Psi_{0}r=0,$ (17)
$(-2rH_{00}-2rH_{0}\Psi_{0}+rH_{0}^{2}+3\Psi_{0}^{2}r+rH_{1}^{2}$
$+\Psi_{1}^{2}r+2r\Psi_{00}-2H_{1}\Psi_{1}r)c_{13}$
$+(rH_{0}^{2}+2r\Psi_{00}-2rH_{0}\Psi_{0}-2rH_{00}+\Psi_{0}^{2}r)c_{2}$
$+(2H_{1}\Psi_{1}r-rH_{1}^{2}-\Psi_{1}^{2}r)c_{3}$
$+(-2H_{1}\Psi_{1}r+rH_{1}^{2}+\Psi_{1}^{2}r)c_{4}$
$+2H_{1}-2\Psi_{1}^{2}r-2\Psi_{0}^{2}r=0,$ (18)
$(3\Psi_{0}^{2}+2H_{1}\Psi_{1}-H_{1}^{2}-\Psi_{1}^{2}+2\Psi_{00}$
$+H_{0}^{2}-2\Psi_{0}H_{0})c_{13}+$
$(H_{0}^{2}+2\Psi_{00}-2\Psi_{0}H_{0}-2H_{00}+\Psi_{0}^{2})c_{2}$
$+(-2H_{1}\Psi_{1}+H_{1}^{2}+\Psi_{1}^{2})c_{3}$
$+(2H_{1}\Psi_{1}-H_{1}^{2}-\Psi_{1}^{2})c_{4}$
$+2\Psi_{1}^{2}-2H_{00}+2H_{11}-2\Psi_{0}^{2}=0,$ (19)
$(-rH_{0}^{2}+2rH_{0}\Psi_{0}+\Psi_{1}^{2}r+2r\Psi_{00}-3\Psi_{0}^{2}r$
$-2H_{1}\Psi_{1}r+rH_{1}^{2})c_{13}+(-rH_{0}^{2}+2rH_{0}\Psi_{0}$
$-\Psi_{0}^{2}r+2rH_{00}-2r\Psi_{00})c_{2}+(2H_{1}\Psi_{1}r$
$-rH_{1}^{2}-\Psi_{1}^{2}r)c_{3}+(-2H_{1}\Psi_{1}r+rH_{1}^{2}$
$+\Psi_{1}^{2}r)c_{4}+4\Psi_{1}-2\Psi_{1}^{2}r+2\Psi_{0}^{2}r-2rH_{11}$
$+4\Psi_{11}r-4r\Psi_{00}+2rH_{00}=0,$ (20)
where the subscript zero denotes the derivative with respect to the coordinate $t$, the subscript one denotes the derivative in relation to the coordinate $r$ and $G^{aether}_{\mu\nu}=G^{Einstein}_{\mu\nu}-T^{aether}_{\mu\nu}$. We assume also $c_{13}=c_{1}+c_{3}$.
4 Gravitational Wave Equation
Obtaining $H_{0}$ from the equation (17), we get
$H_{0}=\frac{\Psi_{0}r[2\Psi_{1}+c_{14}(H_{1}-\Psi_{1})]}{1-c_{14}r(H_{1}-\Psi_{1})},$ (21)
and $H_{00}$ from the equation (20), we have
$H_{00}$ $=$ $\frac{1}{2r(1+c_{2})}\left(-4\Psi_{1}+2\Psi_{1}^{2}r-2\Psi_{0}^{2}r+2c_{14}H_{1}\Psi_{1}r\right.$ (22)
$\left.-2c_{2}rH_{0}\Psi_{0}+c_{13}rH_{0}^{2}+2rH_{11}+4r\Psi_{00}\right.$
$\left.-4\Psi_{11}r-c_{14}rH_{1}^{2}+c_{2}rH_{0}^{2}-c_{14}\Psi_{1}^{2}r+c_{2}\Psi_{0}^{2}r\right.$
$\left.+2c_{2}r\Psi_{00}-2c_{13}rH_{0}\Psi_{0}+3c_{13}\Psi_{0}^{2}r-2c_{13}r\Psi_{00}\right),$
$c_{2}\neq-1$ and $H_{1}$ from the equation (16), we obtain
$H_{1}$ $=$ $\frac{1}{r(c_{13}+c_{2})(-\Psi_{0}+H_{0})}\left(-H_{0}\right.$ (23)
$\left.+2\Psi_{1}\Psi_{0}r+c_{13}H_{0}\Psi_{1}r-3c_{13}\Psi_{1}\Psi_{0}r+c_{2}H_{0}\Psi_{1}r\right.$
$\left.-c_{2}\Psi_{1}\Psi_{0}r+c_{13}H_{0}\right).$
Here, we have checked the possible existence of a bifurcation of solutions. We have two possibilities from equation (23): (i) $c_{2}=c_{13}=0$ or (ii) $c_{2}=-c_{13}$. In general we can assume that $H_{0}\neq\Psi_{0}$. In the case (i) we obtain from the equation (16) and from equation (17) the quantity $H_{0}$ that must be identical to each other in order to have system consistency, i.e., $\Psi_{0}r(-1+2\Psi_{1}r)(-c_{4}+c_{3})(H_{1}-\Psi_{1})=0$. Again, assuming in general that $H_{1}\neq\Psi_{1}$ and substituting the condition $c_{3}=c_{4}$ into equations (15)-(20) we obtain exactly the GR field equations. In the case (ii) we obtain again from the equation (16) and from equation (17) the quantity $H_{0}$ that must be identical to each other in order to have system consistency, i.e., $\Psi_{0}r(-1+2\Psi_{1}r)(c_{14})(H_{1}-\Psi_{1})=0$. Again, assuming in general that $H_{1}\neq\Psi_{1}$ and substituting the condition $c_{14}=0$ into equations (15)-(20) we obtain exactly the EA field equations. These two cases suggest no GR limit for EA wave equations.
We can obtain $H_{11}$ from the equation (19), we compute
$H_{11}$ $=$ $\frac{1}{2}c_{14}H_{1}^{2}+H_{00}-\Psi_{1}^{2}+\Psi_{0}^{2}-\frac{1}{2}c_{2}H_{0}^{2}+$ (24)
$\frac{1}{2}\Psi_{1}^{2}c_{14}-\frac{1}{2}c_{2}\Psi_{0}^{2}+c_{2}H_{00}-c_{2}\Psi_{00}-\Psi_{1}H_{1}c_{14}$
$+c_{2}\Psi_{0}H_{0}-\frac{3}{2}c_{13}\Psi_{0}^{2}-\frac{1}{2}c_{13}H_{0}^{2}-c_{13}\Psi_{00}+c_{13}\Psi_{0}H_{0}.$
Substituting equations (21)-(24) into (19) we obtain the wave equation [26]
${\Psi_{00}}=\frac{1}{1-{c_{13}}}\left(\Psi_{11}+\frac{\Psi_{1}}{r}\right).$ (25)
Obtaining $\Psi_{1}$ from equation (25) we get
$\Psi_{1}=-r(\Psi_{11}-\Psi_{00}+c_{13}\Psi_{00}).$ (26)
Substituting equations (23) and (26) into the equation (16), we obtain an identity.
Substituting equations (21) and (26) into the equation (17) we also obtain an identity.
Substituting equations (21)-(24) and (26) into the equations (19) and (20), we obtain identities.
The components of the field equations that we have not used yet are equations (15) and (18). From equations (15) and (18) we get that
$(-2H_{1}+2\Psi_{1}+2\Psi_{11}r-2rH_{11}-2r\Psi_{00}+2rH_{00})c_{13}$
$+(2rH_{00}-2r\Psi_{00})c_{2}+(2H_{1}-2\Psi_{11}r+2rH_{11}-2\Psi_{1})c_{3}$
$+(2\Psi_{1}+2\Psi_{11}r-2rH_{11}-2H_{1})c_{4}=0.$ (27)
Thus, besides the wave equation (25), we have an additional boundary equation (27) that has to be satisfied in order for gravitational waves to exist in EA theory. Thus, we can calculate also the metric function $H(t,r)$. Notice that in GR where $c_{2}=0$, $c_{3}=0$, $c_{4}=0$ and $c_{13}=0$, we do not have this additional equation. Once we have calculated $\Psi(t,r)$ from the wave equation, we can return to one of the field equations’ components and solve it to obtain $H(t,r)$. Assuming $c_{2}=0$, $c_{3}=0$, $c_{4}=0$, and $c_{13}=0$ in all the calculated results, we obtain exactly the GR results.
5 Solution of the Gravitational Wave Equation
From the equation (25), we can identify that the velocity of the wave is
$c_{ea}={\frac{1}{\sqrt{1-{c_{13}}}}},$ (28)
and the wave number is given by
${k}={\mathrm{w}}\,\sqrt{1-{c_{13}}},$ (29)
where ${\mathrm{w}}$ is the frequency of the gravitational wave. Notice that in EA, the velocity of the wave can be infinity when $c_{13}\rightarrow 1$, while in GR the velocity of the wave is the speed of the light. We also must have $c_{13}<1$ in order to have a wave.
Assuming a stationary wave, we have the monochromatic solution as
$\Psi{(r,t)}=f\left(r\right)\cos\left({\mathrm{w}}\,t\right),$ (30)
and substituting this equation into equation (25) we get
$f\left(r\right)={C_{1}}\,J_{0}\left(k\,r\right)+{C_{2}}\,Y_{0}\left(k\,r\right),$ (31)
where $C_{1}$, $C_{2}$ are arbitrary constants of integration and where $J_{n}(r)$ and $Y_{n}(r)$ are the Bessel of the first and second kind of order $n$, respectively.
Substituting equation (30) and (31) into (27) we obtain that for a monochromatic wave
$H\left(r,t\right)$ $=$ ${F_{1}}\left(r\right){F_{2}}\left(t\right)+\left[{C_{6}}\,J_{0}\left({C_{5}}\,r\right)+{C_{7}}\,Y_{0}\left({C_{5}}\,r\right)\right]\times$ (32)
$\left[{C_{3}}\,{e}^{g(r,t)}+{C_{4}}\,{{e}^{-g(r,t)}}\right],$
where
${F_{1}(r)}={C_{8}}\,J_{0}\left(\sqrt{C_{12}}\,r\right)+{C_{9}}\,Y_{0}\left(\sqrt{C_{12}}\,r\right),$ (33)
${F_{2}(t)}$ $=$ ${C_{10}}\,\cos\left({\frac{\sqrt{{c_{14}}}\,\sqrt{C_{12}}}{\sqrt{{c_{2}}+{c_{13}}}}}\,\,t\right)+{C_{11}}\,\sin\left({\frac{\sqrt{c_{14}}\,\sqrt{C_{12}}}{\sqrt{{c_{2}}+{c_{13}}}}}\,\,t\right),$ (34)
and
$g(r,t)$ $=$ $\sqrt{c_{14}}\times$ (35)
$\left\{\left[-{C_{6}}\left(J_{0}\left({C_{5}}\,r\right)-\frac{J_{1}\left({C_{5}}\,r\right)}{{C_{5}}\,r}\right)-\right.\right.$
$\left.\left.{C_{7}}\left(Y_{0}\left({C_{5}}\,r\right)-\frac{Y_{1}\left({C_{5}}\,r\right)}{{C_{5}}\,r}\right)\right]{{C_{5}}}^{2}\,r-\right.$
$\left.{C_{5}}\,{C_{6}}\,\,J_{1}\left({C_{5}}\,r\right)-{C_{5}}\,{C_{7}}\,\,Y_{1}\left({C_{5}}\,r\right)\right.\Bigg{\}}^{\frac{1}{2}}\times$
${\frac{1}{\sqrt{r}}}\,\,{\frac{1}{\sqrt{{C_{6}}\,J_{0}\left({C_{5}}\,r\right)+{C_{7}}\,Y_{0}\left({C_{5}}\,r\right)}}}\,\,{\frac{1}{\sqrt{{c_{2}}+{c_{13}}}}}\,\,t,$
where $C_{3}$, $C_{4}$, $C_{5}$, $C_{6}$, $C_{7}$, $C_{8}$, $C_{9}$, $C_{10}$, $C_{11}$ and $C_{12}$ are arbitrary constants of integration.
Since the Bessel function of the second kind is singular at $r=0$, we will assume that $C_{2}$, $C_{7}$ and $C_{9}$ are null, thus we have
$f\left(r\right)={C_{1}}\,J_{0}\left(k\,r\right),$ (36)
${F_{1}(r)}={C_{8}}\,J_{0}\left(\sqrt{C_{12}}\,r\right),$ (37)
$H(r,t)$ $=$ ${F_{1}}\left(r\right){F_{2}}\left(t\right)$ (38)
$+{C_{6}}\,{J_{0}\left({C_{5}}\,r\right)}\left[{C_{3}}\,{{e}^{g(r,t)}}+{C_{4}}\,{{e}^{-g(r,t)}}\right].$
Simplifying $g(r,t)$ we get
$g(r,t)\equiv g(t)=i\,\frac{\sqrt{c_{14}}\,\,C_{5}}{\sqrt{c_{13}+c_{2}}}\,t,$ (39)
where $i=\sqrt{-1}$ is the imaginary constant. However, the equation (34) remains unchanged.
Thus, redefining the constants $C_{3}$ and $C_{4}$, we can have
$H(r,t)$ $=$ ${C_{8}}\,{J_{0}\left(\sqrt{C_{12}}\,r\right)}\times$ (40)
$\left[{C_{10}}\,\cos\left({\frac{\sqrt{c_{14}}\,\sqrt{C_{12}}}{\sqrt{{c_{13}}+{c_{2}}}}}\,\,t\right)+{C_{11}}\,\sin\left({\frac{\sqrt{c_{14}}\,\sqrt{C_{12}}}{\sqrt{{c_{13}}+{c_{2}}}}}\,\,t\right)\right]+$
${C_{6}}\,{J_{0}\left({C_{5}}\,r\right)}\times$
$\left[{C_{3}}\,\cos\left({\frac{\sqrt{c_{14}}\,{C_{5}}}{\sqrt{{c_{13}}+{c_{2}}}}}\,\,t\right)+{C_{4}}\,\sin\left({\frac{\sqrt{c_{14}}\,{C_{5}}}{\sqrt{{c_{13}}+{c_{2}}}}}\,\,t\right)\right].$
Substituting these wave solutions into the equations (15)-(20) we must have
${c_{2}}=-{c_{3}}-{{c_{13}}}^{2}-{c_{13}}\,{c_{4}}+{c_{4}}+{c_{13}}\,{c_{3}}.$ (41)
Let us suppose that $\sqrt{C_{12}}=C_{5}=\omega$, $C_{8}=C_{6}$, $C_{11}=C_{4}$, $C_{10}=C_{3}$, for the sake of simplicity, and using equation (29), we can rewrite the metric functions as
$\Psi{(r,t)}={C_{1}}\,\cos\left({\mathrm{w}}\,t\right)\,J_{0}\left(\sqrt{1-{c_{13}}}\,{\mathrm{w}}\,r\right),$ (42)
$H(r,t)$ $=$ $2{C_{8}}\,J_{0}\left(\omega\,r\right)\left[{C_{3}}\,\cos\left(\frac{\sqrt{c_{14}}}{\sqrt{{c_{13}}+{c_{2}}}}\,\omega\,t\right)+{C_{4}}\,\sin\left(\frac{\sqrt{c_{14}}}{\sqrt{{c_{13}}+{c_{2}}}}\,\omega\,t\right)\right],$
Let us suppose that $C_{8}=1/2$, $C_{3}=1$, $C_{4}=0$ for the sake of simplicity. Thus, we can rewrite the metric functions as
$H(r,t)$ $=$ $\cos\left(\frac{\sqrt{c_{14}}}{\sqrt{{c_{13}}+{c_{2}}}}\,\omega\,\,t\right)J_{0}\left(\omega\,r\right).$ (44)
Comparing equations (42) and (44), we can see that they have the same functional dependence in the coordinates $r$ and $t$, but with different frequencies in both coordinates, ${\mathrm{w}}$ and $\omega$. Section 6 will show that this temporal behavior is different in GR.
6 Gravitational Wave Pulse with Width
Let us now calculate the superposition of waves of the type of (42) and (44) with amplitude factor as $\exp(-\beta\,{\mathrm{w}})$. Here, we will follow the same procedure used in reference [34]. Thus, we have
[…]
From the reference [35] we have that
[…]
$\frac{\sqrt{2}}{2}\,{\frac{\sqrt{\sqrt{{{\beta}}^{4}+2\,{{\beta}}^{2}{{\gamma}}^{2}+2\,{{\alpha}}^{2}{{\beta}}^{2}+{{\gamma}}^{4}-2\,{{\gamma}}^{2}{{\alpha}}^{2}+{{\alpha}}^{4}}+{{\beta}}^{2}+{{\gamma}}^{2}-{{\alpha}}^{2}}}{\sqrt{{{\beta}}^{4}+2\,{{\beta}}^{2}{{\gamma}}^{2}+2\,{{\alpha}}^{2}{{\beta}}^{2}+{{\gamma}}^{4}-2\,{{\gamma}}^{2}{{\alpha}}^{2}+{{\alpha}}^{4}}}},$ (46)
$\Psi(r,t)=$
$\frac{\sqrt{2}}{2}{\frac{\sqrt{\sqrt{{{\beta}}^{4}+2\,{{\beta}}^{2}{r}^{2}+2\,{t}^{2}{{\beta}}^{2}+{r}^{4}-2\,{r}^{2}{t}^{2}+{t}^{4}}+{{\beta}}^{2}+{r}^{2}-{t}^{2}}}{\sqrt{{{\beta}}^{4}+2\,{{\beta}}^{2}{r}^{2}+2\,{t}^{2}{{\beta}}^{2}+{r}^{4}-2\,{r}^{2}{t}^{2}+{t}^{4}}}},$ (47)
where $\alpha=t$ and $\gamma=\sqrt{1-c_{13}}\,r$.
Let us make the following coordinate transformation
$r=-t+\beta\,x$ (48)
assuming that $\beta=1$, $t=0$ or $t=2$ or $t=6$ or $t=10$ and $c_{13}=0$ or $c_{13}=\pm 1/2$, we can write that
$\Psi(x,t=0,c_{13}=1/2)$ $=$ $\,{\frac{\sqrt{2}}{\sqrt{2+{x}^{2}}}},$ (49)
$\Psi(x,t=2,c_{13}=1/2)=$
$\frac{1}{2}\,{\frac{\sqrt{2\,\sqrt{100-12\,\left(-2+x\right)^{2}+\left(-2+x\right)^{4}}-12+2\,\left(-2+x\right)^{2}}\sqrt{2}}{\sqrt{100-12\,\left(-2+x\right)^{2}+\left(-2+x\right)^{4}}}},$ (50)
$\Psi(x,t=6,c_{13}=1/2)=$
$\frac{1}{2}\,{\frac{\sqrt{2\,\sqrt{5476-140\,\left(-6+x\right)^{2}+\left(-6+x\right)^{4}}-140+2\,\left(-6+x\right)^{2}}\sqrt{2}}{\sqrt{5476-140\,\left(-6+x\right)^{2}+\left(-6+x\right)^{4}}}},$ (51)
$\Psi(x,t=10,c_{13}=1/2)=$
$\frac{1}{2}\,{\frac{\sqrt{2\,\sqrt{40804-396\,\left(-10+x\right)^{2}+\left(-10+x\right)^{4}}-396+2\,\left(-10+x\right)^{2}}\sqrt{2}}{\sqrt{40804-396\,\left(-10+x\right)^{2}+\left(-10+x\right)^{4}}}},$
$\Psi(x,t=0,c_{13}=0)$ $=$ ${\frac{1}{\sqrt{1+{x}^{2}}}},$ (53)
$\Psi(x,t=2,c_{13}=0)=$
$\frac{1}{2}\,{\frac{\sqrt{\sqrt{25-6\,\left(-2+x\right)^{2}+\left(-2+x\right)^{4}}-3+\left(-2+x\right)^{2}}\sqrt{2}}{\sqrt{25-6\,\left(-2+x\right)^{2}+\left(-2+x\right)^{4}}}},$ (54)
$\Psi(x,t=6,c_{13}=0)=$
$\frac{1}{2}\,{\frac{\sqrt{\sqrt{1369-70\,\left(-6+x\right)^{2}+\left(-6+x\right)^{4}}-35+\left(-6+x\right)^{2}}\sqrt{2}}{\sqrt{1369-70\,\left(-6+x\right)^{2}+\left(-6+x\right)^{4}}}},$ (55)
$\Psi(x,t=10,c_{13}=0)=$
$\frac{1}{2}\,{\frac{\sqrt{\sqrt{10201-198\,\left(-10+x\right)^{2}+\left(-10+x\right)^{4}}-99+\left(-10+x\right)^{2}}\sqrt{2}}{\sqrt{10201-198\,\left(-10+x\right)^{2}+\left(-10+x\right)^{4}}}}.$ (56)
$\Psi(x,t=0,c_{13}=-1/2)$ $=$ ${\frac{\sqrt{2}}{\sqrt{2+3\,{x}^{2}}}},$ (57)
$\Psi(x,t=2,c_{13}=-1/2)=$
$\frac{1}{2}\,{\frac{\sqrt{2\,\sqrt{100-36\,\left(-2+x\right)^{2}+9\,\left(-2+x\right)^{4}}-12+6\,\left(-2+x\right)^{2}}\sqrt{2}}{\sqrt{100-36\,\left(-2+x\right)^{2}+9\,\left(-2+x\right)^{4}}}},$ (58)
$\Psi(x,t=6,c_{13}=-1/2)=$
$\frac{1}{2}\,{\frac{\sqrt{2\,\sqrt{5476-420\,\left(-6+x\right)^{2}+9\,\left(-6+x\right)^{4}}-140+6\,\left(-6+x\right)^{2}}\sqrt{2}}{\sqrt{5476-420\,\left(-6+x\right)^{2}+9\,\left(-6+x\right)^{4}}}},$
$\Psi(x,t=10,c_{13}=-1/2)=$
$\frac{1}{2}\,{\frac{\sqrt{2\,\sqrt{40804-1188\,\left(-10+x\right)^{2}+9\,\left(-10+x\right)^{4}}-396+6\,\left(-10+x\right)^{2}}\sqrt{2}}{\sqrt{40804-1188\,\left(-10+x\right)^{2}+9\,\left(-10+x\right)^{4}}}}.$
See these examples in Figures 2 to 7, using the parameters of the Table 1.
Table 1: Summary of the Aether Parameters
Case $c_{13}$ $c_{14}$ $c_{2}$
$A$1/2 2 7/2
$B$ 0 1/2 1/2
$C$ 1/2 3/2 1/4
These parameters must obey the equation (41) and the solar system tests, thus, we must also have $20\leq c_{14}<$ 2, $c_{13}<1$ and $02+c_{13}+2c_{2}>$ 0 [36].
Figure 1: Wave components $\Psi({\mathrm{w}})$ and $H(\omega)$ using $r=1$, $t=1$ and the parameters of Table 1 in equations (42) and (44).
Figure 2: Wave component pulse $\Psi$ of width $\beta$, using $C_{1}=1$ in equation (42) and $\alpha=t$, $\beta=1$, $\gamma=\sqrt{1-c_{13}}\,r$ in equation (46), for four values of $t=0$, $t=2$, $t=6$ and $t=10$ and the parameters of Table 1.
Figure 3: Wave component pulse $\Psi$ of width $\beta$, using $C_{1}=1$ in equation (42) and $\alpha=t$, $\beta=1$, $\gamma=\sqrt{1-c_{13}}\,r$ in equation (46), for four values of $t=0$, $t=2$, $t=6$ and $t=10$, and the parameters of Table 1.
Figure 4: Wave component pulse $\Psi$ of width $\beta$, using $C_{1}=1$ in equation (42) and $\alpha=t$, $\beta=1$, $\gamma=\sqrt{1-c_{13}}\,r$ in equation (46), for four values of $t=0$, $t=2$, $t=6$ and $t=10$ and the parameters of Table 1.
Figure 5: Wave component pulse $H$ of width $\beta$, using equation (44) and $\alpha={t\,\sqrt{c_{14}}}/{\sqrt{{c_{13}}+{c_{2}}}}$, $\beta=1$, $\gamma=r$ in equation (46), for four values of $t=0$, $t=2$, $t=6$ and $t=10$ and the parameters of Table 1.
Figure 6: Wave component pulse $H$ of width $\beta$, using equation (44) and $\alpha={t\,\sqrt{c_{14}}}/{\sqrt{{c_{13}}+{c_{2}}}}$, $\beta=1$, $\gamma=r$ in equation (46), for four values of $t=0$, $t=2$, $t=6$ and $t=10$, and the parameters of Table 1.
Figure 7: Wave component pulse $H$ of width $\beta$, using in equation (44) and $\alpha={t\,\sqrt{c_{14}}}/{\sqrt{{c_{13}}+{c_{2}}}}$, $\beta=1$, $\gamma=r$ in equation (46), for four values of $t=0$, $t=2$, $t=6$ and $t=10$ and the parameters of Table 1.
7 Comparison of $\bf\Psi(r,t)$ and $\bf H(r,t)$ in EA and GR
Assuming (13) and $c_{1}=c_{2}=c_{3}=c_{4}=0$ we compute the different terms in the field equations (12), giving
$G_{00}$ $=$ $2H_{1}-2\Psi_{1}^{2}r-2\Psi_{0}^{2}r=0,$ (61)
$G_{01}$ $=$ $H_{0}+2\Psi_{1}\Psi_{0}r=0,$ (62)
$G_{11}$ $=$ $2H_{1}-2\Psi_{1}^{2}r-2\Psi_{0}^{2}r=0,$ (63)
$G_{22}$ $=$ $2\Psi_{1}^{2}-2H_{00}+2H_{11}-2\Psi_{0}^{2}=0,$ (64)
$G_{33}$ $=$ $0\displaystyle 4\Psi_{1}-2\Psi_{1}^{2}r+2\Psi_{0}^{2}r-2rH_{11}+4\Psi_{11}r-4r\Psi_{00}+2rH_{00}=$ 0 . (65)
The wave equation for GR is obtained using (65) and (64), thus
${\Psi_{00}}=\left(\Psi_{11}+\frac{\Psi_{1}}{r}\right).$ (66)
Substituting $\Psi_{1}$ from equation (66) into equations (62) and (61) we get
${H_{0}}-2\,{\Psi_{0}}\,{r}^{2}{\Psi_{00}}+2\,{\Psi_{0}}\,{r}^{2}{\Psi_{11}}=0,$ (67)
${H_{1}}-{r}^{3}{{\Psi_{00}}}^{2}+2\,{r}^{3}{\Psi_{00}}\,{\Psi_{11}}-{{\Psi_{11}}}^{2}{r}^{3}-{{\Psi_{0}}}^{2}r=0,$ (68)
This must be solved to obtain $H(r,t)$. Notice that these two equations are not equivalent to the equation (27) in EA. In GR these equations come from the components $G_{01}$ and $G_{00}$, while in EA the equation (27) comes from equations (18) and (15).
Solving equation (66) we get
$\Psi(r,t)=\left[{C_{13}}\,{J_{0}\left({\mathrm{w}}r\right)}+{C_{14}}\,{Y_{0}\left({\mathrm{w}}r\right)}\right]\cos\left({\mathrm{w}}t\right).$ (69)
We will assume $C_{14}=0$ since $Y_{0}$ is singular at $r=0$ and $C_{13}=1$ for simplicity. Thus,
$\Psi(r,t)={J_{0}\left({\mathrm{w}}r\right)}\cos\left({\mathrm{w}}t\right).$ (70)
Substituting $\Psi$ from (69) into (67) and (68) and solving both equations, we obtain
$H(r,t)$ $=$ $-\frac{1}{2}\,{\mathrm{w}}r\left\{2\,\cos\left({\mathrm{w}}t\right)^{2}J_{0}\left({\mathrm{w}}r\right)J_{1}\left({\mathrm{w}}r\right)-\right.$ (71)
$\left.{\mathrm{w}}r\left[J_{0}\left({\mathrm{w}}r\right)^{2}+J_{1}\left({\mathrm{w}}r\right)^{2}\right]\right\}.$
See Figure 8 for $\Psi({\mathrm{w}})$ and $H(\omega)$.
Notice that the last term of this equation is not periodic in time, turning the solution of the gravitational wave into a quasi-periodic one. This differs entirely from the EA gravitational wave, where the metric functions $\Psi(r,t)$ and $H(r,t)$ are both completely periodic in time.
Substituting $\cos\left({\mathrm{w}}t\right)^{2}=[1+\cos\left(2{\mathrm{w}}t\right)]/2$ and $\cos\left(2{\mathrm{w}}t\right)=[\exp\left(2i{\mathrm{w}}t\right)+\exp\left(-2i{\mathrm{w}}t\right)]/2$ into equation (71), we can integrate it and obtain it in terms of the Meijer G function [37], thus
$H(r,t)$ $=$ $-\frac{1}{2\beta\pi}\,G^{2,1}_{2,2}\left({\frac{{\beta}^{2}}{4{r}^{2}}}\,\Big{|}\,^{0,1}_{\frac{1}{2},\frac{1}{2}}\right)-$ (72)
$\frac{1}{4\pi\left(-2\,it+\beta\right)}\,G^{2,1}_{2,2}\left({\frac{\left(-2\,it+\beta\right)^{2}}{4{r}^{2}}}\,\Big{|}\,^{0,1}_{\frac{1}{2},\frac{1}{2}}\right)-$
$\frac{1}{4\pi\left(\beta+2\,it\right)}\,G^{2,1}_{2,2}\left({\frac{\left(\beta+2\,it\right)^{2}}{4{r}^{2}}}\,\Big{|}\,^{0,1}_{\frac{1}{2},\frac{1}{2}}\right)-$
$\frac{r}{{\beta}^{2}\pi}\left[G^{2,2}_{3,3}\left({\frac{{\beta}^{2}}{4{r}^{2}}}\,\Big{|}\,^{\frac{1}{2},\frac{1}{2},\frac{1}{2}}_{1,0,\frac{3}{2}}\right)-G^{2,1}_{2,2}\left({\frac{{\beta}^{2}}{4{r}^{2}}}\,\Big{|}\,^{-\frac{1}{2},\frac{1}{2}}_{1,0}\right)\right]$
See Figures 9 and 10 for the pulse in GR.
Figure 8: Wave components $\Psi({\mathrm{w}})$ and $H(\omega)$ using $r=1$ and $t=1$ in equations (70) and (71).
Figure 9: Wave component pulse $H$ of width $\beta$, using equation (44) and $\alpha=t$, $\beta=1$, $\gamma=r$ in equation (46), for four values of $t=0$, $t=2$, $t=6$ and $t=10$. Fitting a straight line, we get $H_{s}\approx 0.1862+0.3186\,x$.
Figure 10: Wave component pulse $H$ of width $\beta$, using equation (44) and $\alpha=t$, $\beta=1$, $\gamma=r$ in equation (46), for four values of $t=0$, $t=2$, $t=6$ and $t=10$.
8 Energy and Momentum of Cylindrical Gravitational Waves
The gravitational energy-momentum pseudotensor, which Einstein initially formulated in 1916 following his discovery of the general relativity field equations, closely connects with the conventional canonical stress tensor applicable to matter fields in a flat spacetime. Many years ago, Scheidegger [38] questioned the well-defined existence of gravitational radiation. In response, Rosen [39] explored the presence of energy and momentum in cylindrical gravitational waves. Employing the energy-momentum pseudotensors of Einstein and those of Landau and Lifshitz, Rosen calculated using cylindrical polar coordinates and determined that energy and momentum density components were null. These findings aligned with Scheidegger’s suggestion that a physical system could not emit gravitational energy. However, Rosen [40] later acknowledged an error and re-evaluated the calculations in cartesian coordinates, revealing nonvanishing and reasonable energy and momentum density. Despite Rosen’s initial promise to publish the details elsewhere, this was not accomplished until later when he and another author [41] re-examined the energy-momentum density components in cartesian coordinates following Einstein’s prescription. Their subsequent findings were finite and reasonable [42]. Here, we compute these quantities in EA theory using the equivalent expressions for the EA theory derived in reference [43].
The Einstein pseudotensor and the von Freud superpotential [44] obey the following relations
$\partial_{\mu}t_{\nu}{}^{\mu}$ $=$ $0$ (73)
$t_{\nu}{}^{\mu}$ $=$ $-\partial_{\gamma}U_{\nu}{}^{\gamma\mu}.$ (74)
We present the pure Einstein part (GR) and pure aether contributions separately for easy comparison.
${}^{\rm Einstein}t^{\ \mu}_{\rho}$ $=$ $\frac{\sqrt{-g}}{16\pi G}\left(\delta^{\ \mu}_{\rho}(\Gamma^{\alpha}_{\beta\gamma}\Gamma^{\beta}_{\alpha\delta}-\Gamma^{\alpha}_{\alpha\beta}\Gamma^{\beta}_{\gamma\delta})g^{\gamma\delta}\right.$ (75)
$\left.+\Gamma^{\beta}_{\rho\alpha}\Gamma^{\gamma}_{\gamma\beta}g^{\mu\alpha}-\Gamma^{\beta}_{\rho\beta}\Gamma^{\gamma}_{\gamma\alpha}g^{\mu\alpha}+\Gamma^{\alpha}_{\rho\alpha}\Gamma^{\mu}_{\beta\gamma}g^{\beta\gamma}\right.$
$\left.+\Gamma^{\mu}_{\rho\alpha}\Gamma^{\beta}_{\beta\gamma}g^{\alpha\gamma}-2\Gamma^{\mu}_{\alpha\beta}\Gamma^{\alpha}_{\rho\gamma}g^{\beta\gamma}\right),$
${}^{\rm Aether}t_{\nu}{}^{\lambda}$ $=$ $\frac{1}{16\pi G}\left(2{\sqrt{-g}}J^{\lambda}{}_{\rho}\nabla_{\nu}u^{\rho}-\sqrt{-g}\{(J^{\lambda}{}_{\beta}+J_{\beta}{}^{\lambda})u^{\alpha}\right.$ (76)
$\left.-(J^{\alpha}{}_{\beta}+J_{\beta}{}^{\alpha})u^{\lambda}+(J^{\lambda\alpha}-J^{\alpha\lambda})u_{\beta}\}\Gamma^{\beta}{}_{\alpha\nu}\right.$
$\left.+\delta_{\nu}^{\lambda}{\sqrt{-g}}L_{u}\right),$
${}^{\rm Einstein}U_{\beta}{}^{\lambda\alpha}=\frac{1}{16\pi G}\frac{1}{\sqrt{-g}}g_{\beta\tau}\partial_{\gamma}\{(-g)(g^{\lambda\tau}g^{\alpha\gamma}-g^{\alpha\tau}g^{\lambda\gamma})\},$ (77)
and
${}^{\rm Aether}U_{\beta}{}^{\lambda\alpha}$ $=$ $\frac{1}{16\pi G}\sqrt{-g}\left((J^{\lambda}{}_{\beta}+J_{\beta}{}^{\lambda})u^{\alpha}\right.$ (78)
$\left.-(J^{\alpha}{}_{\beta}+J_{\beta}{}^{\alpha})u^{\lambda}+(J^{\lambda\alpha}-J^{\alpha\lambda})u_{\beta}\right),$
where $J^{a}{}_{\beta}$ is defined in (10). The above decompositions of $t_{\nu}{}^{\mu}$ and $U_{\nu}{}^{\gamma\mu}$ into pure Einstein and aether pieces do not satisfy (73) and (74) independently. When evaluating these pseudotensorial expressions, the metric must have connection coefficients vanish in $O(1/r)$ or faster in the asymptotic limit in that coordinate system. If the coordinates are not chosen properly, these expressions will yield incorrect energies and momenta. Further details on this can be found in [43].
Let us now calculate the energy-momentum pseudotensors given by the equations (75) and (76) in cartesian coordinates, i.e.,
$ds^{2}=$ ${\frac{{e}^{-2\,\psi}\left({e}^{2\,H}{x}^{2}+{y}^{2}\right)\,dx^{2}}{{x}^{2}+{y}^{2}}}+2\,{\frac{{e}^{-2\,\psi}\left({e}^{2\,H}-1\right)xy\,dx\,dy}{{x}^{2}+{y}^{2}}}+$ (79)
${\frac{{e}^{-2\,\psi}\left({e}^{2\,H}{y}^{2}+{x}^{2}\right)\,dy^{2}}{{x}^{2}+{y}^{2}}}+{e}^{2\,\psi}\,dz^{2}-{e}^{2\,H-2\,\psi}\,d{t}^{2}.$
In order to make the coordinate transformation from the cylindrical to the cartesian coordinates, i.e., $\bar{\psi}(r,\theta,t)\rightarrow\psi\left(x,y,t\right)$ and $\bar{H}(r,\theta,t)\rightarrow H\left(x,y,t\right)$, let us use the usual polar coordinate transformation that says $x(r,\theta)=r\cos(\theta)$ and $y(r,\theta)=r\sin(\theta)$. 111 Let us define now an arbitrary function $F$ in polar coordinates given by $\bar{F}\left(r,\theta\right)=F\left(x(r,\theta),\,y(r,\theta)\right)$. The chain rule says that $\bar{F}_{r}=F_{x}\,x_{r}+F_{y}\,y_{r}$ and $\bar{F}_{\theta}=F_{x}\,x_{\theta}+F_{y}\,y_{\theta}$. In our case, we must impose that $\bar{F}_{\theta}=0$. Thus, we will be able to find the correct coordinate transformations, i.e., $F_{x}=\bar{F}_{r}\cos(\theta)$; $F_{y}=\bar{F}_{r}\sin(\theta)$; $F_{xx}=\bar{F}_{rr}\cos(\theta)^{2}$; $F_{yy}=\bar{F}_{rr}\sin(\theta)^{2}$; $F_{xy}=\bar{F}_{rr}\cos(\theta)\sin(\theta)$; $F_{tx}=\bar{F}_{tr}\cos(\theta)$ and $F_{ty}=\bar{F}_{tr}\sin(\theta)$.
The components of the energy-momentum pseudotensor are given by the equation (75) are
${}^{Einstein}t_{x}{}^{x}$ $=$ $\frac{1}{8\pi G}\,\left[{\frac{\left(x-y\right)\left(y+x\right){\psi_{{r}}}^{2}}{{r}^{2}}}+{{{\psi_{{t}}}^{2}}}\right],$ (80)
${}^{Einstein}t_{y}{}^{x}$ $=$ ${\frac{1}{4\pi G}\,\frac{{\psi_{{r}}}^{2}xy}{{r}^{2}}},$ (81)
${}^{Einstein}t_{x}{}^{y}={\frac{1}{4\pi G}\,\frac{{\psi_{{r}}}^{2}xy}{{r}^{2}}},$ (82)
${}^{Einstein}t_{y}{}^{y}$ $=$ $-\frac{1}{8\pi G}\,\left[{\frac{\left(x-y\right)\left(y+x\right){\psi_{{r}}}^{2}}{{r}^{2}}-{{\psi_{{t}}}^{2}}}\right],$ (83)
${}^{Einstein}t_{z}{}^{z}$ $=$ $-\frac{1}{8\pi G}\,\left[{{e}^{2\,H}{\psi_{{r}}}^{2}+{e}^{2\,H}{\psi_{{t}}}^{2}-2\,{\psi_{{t}}}^{2}}\right],$ (84)
${}^{Einstein}t_{t}{}^{t}$ $=$ $-\frac{1}{8{\pi G}}\,{\left({\psi_{{r}}}^{2}+{\psi_{{t}}}^{2}\right){e}^{2\,H}},$ (85)
${}^{Einstein}t_{x}{}^{t}=-\frac{1}{4\pi G}\,{\frac{x\psi_{{r}}\psi_{{t}}}{r}},$ (86)
${}^{Einstein}t_{t}{}^{x}$ $=$ $\frac{1}{4\pi G}\,{\frac{x\psi_{{r}}\psi_{{t}}{e}^{2\,H}}{r}},$ (87)
${}^{Einstein}t_{y}{}^{t}$ $=$ $-\frac{1}{4\pi G}\,{\frac{y\psi_{{r}}\psi_{{t}}}{r}},$ (88)
${}^{Einstein}t_{t}{}^{y}$ $=$ $\frac{1}{4\pi G}\,{\frac{y\psi_{{r}}\psi_{{t}}{e}^{2\,H}}{r}},$ (89)
${}^{Einstein}t_{z}{}^{t}$ $=$ ${}^{Einstein}t_{t}{}^{z}=0,$ (90)
where the subscript notation means $F_{v}=\frac{\partial F}{\partial v}$ ($F$ is an arbitrary function $F=F(r,t)$).
We can compare the components (85)-(90) with the papers [41] and [42]. We can notice that they are the same unless the minus signal in all our components due to the different signature of the metric assumed in this work.
The components of the energy-momentum pseudotensor are given by the equation (76) are
${}^{Aether}t_{x}{}^{x}=-\frac{1}{16\pi\,G}$ $\left[{\frac{\left({r}^{2}{c_{1}}+{r}^{2}{c_{4}}-2\,{c_{3}}\,{x}^{2}\right){H_{{r}}}^{2}}{{r}^{2}}}-\right.$ (91)
$\left.{\frac{2\left({r}^{2}{c_{1}}+{r}^{2}{c_{4}}-2\,{c_{3}}\,{x}^{2}\right)\psi_{{r}}H_{{r}}}{{r}^{2}}}-\right.$
$\left.{\left({c_{1}}+{c_{2}}+{c_{3}}\right){H_{{t}}}^{2}}+{{2\left({c_{1}}+{c_{2}}+{c_{3}}\right)\psi_{{t}}H_{{t}}}}+\right.$
$\left.{\frac{\left({r}^{2}{c_{1}}+{r}^{2}{c_{4}}-2\,{c_{3}}\,{x}^{2}\right){\psi_{{r}}}^{2}}{{r}^{2}}}-{\left(3\,{c_{1}}+{c_{2}}+3\,{c_{3}}\right){\psi_{{t}}}^{2}}\right],$
${}^{Aether}t_{y}{}^{x}$ $=$ $\frac{1}{8\pi G}\,{\frac{y{c_{3}}\,x\left(H_{{r}}-\psi_{{r}}\right)^{2}}{{r}^{2}}},$ (92)
${}^{Aether}t_{x}{}^{y}=\frac{1}{8\pi G}\,{\frac{y{c_{3}}\,x\left(H_{{r}}-\psi_{{r}}\right)^{2}}{{r}^{2}}},$ (93)
${}^{Aether}t_{y}{}^{y}=-\frac{1}{16{\pi G}}$ $\left[{\frac{\left({r}^{2}{c_{1}}+{r}^{2}{c_{4}}-2\,{c_{3}}\,{y}^{2}\right){H_{{r}}}^{2}}{{r}^{2}}}-\right.$ (94)
$\left.{\frac{2\left({r}^{2}{c_{1}}+{r}^{2}{c_{4}}-2\,{c_{3}}\,{y}^{2}\right)\psi_{{r}}H_{{r}}}{{r}^{2}}}-\right.$
$\left.{\left({c_{1}}+{c_{2}}+{c_{3}}\right){H_{{t}}}^{2}}+{2\left({c_{1}}+{c_{2}}+{c_{3}}\right)\psi_{{t}}H_{{t}}}+\right.$
$\left.{\frac{\left({r}^{2}{c_{1}}+{r}^{2}{c_{4}}-2\,{c_{3}}\,{y}^{2}\right){\psi_{{r}}}^{2}}{{r}^{2}}}-{\left(3\,{c_{1}}+{c_{2}}+3\,{c_{3}}\right){\psi_{{t}}}^{2}}\right],$
${}^{Aether}t_{z}{}^{z}=-\frac{1}{16\pi G}$ $\left[{\left({c_{1}}+{c_{4}}\right){H_{{r}}}^{2}}-{2\left({c_{1}}+{c_{4}}\right)\psi_{{r}}H_{{r}}}-{\left({c_{1}}+{c_{2}}+{c_{3}}\right){H_{{t}}}^{2}}+\right.$
$\left.{2\left({c_{1}}+{c_{2}}+{c_{3}}\right)\psi_{{t}}H_{{t}}}+{\left({c_{1}}+{c_{4}}\right){\psi_{{r}}}^{2}}-{\left(3\,{c_{1}}+{c_{2}}+3\,{c_{3}}\right){\psi_{{t}}}^{2}}\right],$
${}^{Aether}t_{t}{}^{t}=-\frac{1}{16\pi G}$ $\left[{\left(3\,{H_{{r}}}^{2}-6\,\psi_{{r}}H_{{r}}+3\,{\psi_{{r}}}^{2}+3\,{\psi_{{t}}}^{2}-2\,H_{{t}}\psi_{{t}}+{H_{{t}}}^{2}\right){c_{1}}}+\right.$ (96)
$\left.{\left(H_{{t}}-\psi_{{t}}\right)^{2}{c_{2}}}-\right.$
$\left.{\left(2\,{H_{{r}}}^{2}-4\,\psi_{{r}}H_{{r}}+2\,{\psi_{{r}}}^{2}-3\,{\psi_{{t}}}^{2}+2\,H_{{t}}\psi_{{t}}-{H_{{t}}}^{2}\right){c_{3}}}+\right.$
$\left.{3\left(H_{{r}}-\psi_{{r}}\right)^{2}{c_{4}}}\right],$
${}^{Aether}t_{x}{}^{t}=-\frac{1}{8\pi G}\,$ $\left[{\frac{x\psi_{{t}}\left(2\,r\psi_{{r}}+{e}^{2\,H}-1\right){c_{1}}}{{r}^{2}}}-\right.$ (97)
$\left.\frac{x\left(H_{{t}}-\psi_{{t}}\right)\left(-r\psi_{{r}}+H_{{r}}r+{e}^{2\,H}-1\right){c_{2}}}{{r}^{2}}+\right.$
$\left.{\frac{x\psi_{{t}}\left(2\,r\psi_{{r}}+{e}^{2\,H}-1\right){c_{3}}}{{r}^{2}}}+\right.$
$\left.{\frac{x\left(H_{{t}}-\psi_{{t}}\right)\left(H_{{r}}-\psi_{{r}}\right){c_{4}}}{r}}\right],$
${}^{Aether}t_{t}{}^{x}=-\frac{1}{8\pi G}$ $\left[{\frac{x\left(H_{{t}}-\psi_{{t}}\right)\left(H_{{r}}-\psi_{{r}}\right)\left(2\,{c_{1}}-2\,{c_{2}}-{c_{3}}\right)}{r}}\right],$ (98)
${}^{Aether}t_{y}{}^{t}=\frac{1}{8\pi G}$ $\left[-{\frac{y\psi_{{t}}\left(2\,r\psi_{{r}}+{e}^{2\,H}-1\right){c_{1}}}{{r}^{2}}}+\right.$ (99)
$\left.{\frac{y\left(H_{{t}}-\psi_{{t}}\right)\left(-r\psi_{{r}}+H_{{r}}r+{e}^{2\,H}-1\right){c_{2}}}{{r}^{2}}}-\right.$
$\left.\frac{y\psi_{{t}}\left(2\,r\psi_{{r}}+{e}^{2\,H}-1\right){c_{3}}}{{r}^{2}}-\right.$
$\left.{\frac{y\left(H_{{t}}-\psi_{{t}}\right)\left(H_{{r}}-\psi_{{r}}\right){c_{4}}}{r}}\right],$
${}^{Aether}t_{t}{}^{y}$ $=$ $\frac{1}{8\pi G}\,{\frac{y\left(2{c_{1}}+2{c_{2}}+{c_{3}}\right)\left(H_{{t}}H_{{r}}-\psi_{{t}}H_{{r}}-H_{{t}}\psi_{{r}}+\psi_{{t}}\psi_{{r}}\right)}{r}},$ (100)
${}^{Aether}t_{z}{}^{t}$ $=$ ${}^{Aether}t_{t}{}^{z}=0,$ (101)
As said before, when evaluating the energy-momentum pseudotensor, the metric must be written in a coordinate system where the connection coefficients vanish like $\mathcal{O}(1/r)$ or faster in the asymptotic limit. Thus, we can see from equations (80)-(101) that the energy and momentum densities are finite and reasonable at infinity. From equations (42) and (44), we can notice that the quantities $H_{t}$ and $\psi_{t}$ are oscillatory functions of the time. Since the quantities $H_{r}$ and $\psi_{r}$ are also oscillatory in the radial coordinate and decreasing at infinity. Thus, we have finite values of the energy-momentum pseudotensors at infinity.
We can notice that EA energy-momentum pseudotensors are completely different from the GR ones by construction, i.e., if we assume $c_{1}=c_{2}=c_{3}=c_{4}=$ all the EA energy-momentum pseudotensor components are zero.
9 Conclusions
We have obtained the gravitational wave equation for cylindrical symmetry in the EA theory. This is the generalization of the Einstein-Rosen wave equation in GR. We have shown that the gravitational wave in the EA is periodic in time, both for $\Psi(r,t)$ and $H(r,t)$ metric functions. However, in GR, $\Psi(r,t)$ is periodic in time, but $H(r,t)$ is semi-periodic in time, having a secular drifting in the wave frequency. Due to this frequency drifting, the evolution of wave pulses of a given width emitted is entirely different in both theories in the $H(r,t)$ metric function. Another difference between the two theories is the gravitational wave velocity. While in GR, the waves propagate with the speed of light, in EA, there is not an upper limit in the wave velocity, reaching infinity value if $c_{13}\rightarrow 1$ and zero if $c_{13}\rightarrow-\infty$. Also, aether has contributions to energy-momentum pseudotensor and superpotential. In principle, all these characteristics could be a way to differentiate GR and EA theories observationally.
10 Acknowledgments
The author (RC) acknowledges the financial support from FAPERJ (E-26/171.754/2000, E-26/171.533/2002 and E-26/170.951/2006). MFAdaS acknowledges the financial support from CNPq-Brazil, FINEP-Brazil (Ref. 2399/03), FAPERJ/UERJ (307935/2018-3), FAPERJ (E-26/211.906/2021) and from CAPES (CAPES-PRINT 41/2017).